# Simulating non-Markovian open quantum dynamics by exploiting physics-informed neural network

### 09 Mar 2026 v1

@@ III.1 Simulation of non-Markovian dissipative dynamics of Anderson impurity model
We consider an OQS comprising a localized impurity symmetrically coupled to two noninteracting electron reservoirs. This setup is important for understanding electron transport through molecular junctions. The OQS is described by the single-impurity Anderson model Anderson (1961). The impurity Hamiltonian is $H_{{}_{\rm S}}(t)=\epsilon_{0}(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+U_{0}\hat{n}_{\uparrow}\hat{n}_{\downarrow}+{\Theta}(t-t_{0})[\Delta\epsilon(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+\Delta U\hat{n}_{\uparrow}\hat{n}_{\downarrow}$], where $\hat{n}_{s}$ is the occupation operator for spin-$s$ electrons, $\epsilon_{0}$ is the impurity energy, and $U_{0}$ is the Coulomb interaction energy. At time $t_{0}=0$, a sudden change shifts the impurity’s parameters by $\Delta\epsilon$ and $\Delta U$, and a bias voltage is simultaneously switched on, establishing a chemical potential difference between the reservoirs. The inset of Fig. 3(a) depicts the resulting open quantum dynamics. The impurity level shift triggers electron transfer from the reservoirs to the impurity on a relatively short timescale. Eventually a steady electric current flows through the impurity. At sufficiently low reservoir temperatures, environmental memory effects influence the electron transition process, resulting in non-Markovian oscillations. In the following, we also calculate the non-Markovian open quantum dynamics using the fermionic HEOM method implemented in the HEOM-QUICK program Ye et al. (2016); Zhang et al. (2024), and the numerical outcomes are served as reference values for assessing the accuracy of the PINN-DQME method.
@@
Figure 3: (a) The evolution of the electric current flowing into the impurity from the right reservoir at different temperatures. Inset: schematic of quantum dissipative dynamics for an impurity coupled to left (L) and right (R) reservoirs with chemical potentials $\mu_{\rm L}$ and $\mu_{\rm R}$. (b) The evolution of the occupation number of the up spin electron in the impurity at two different temperatures. The purple and red dots represent the decomposition points $\{t_{p}\}$ for the time domain. The results of PINN-DQME (solid lines) are benchmarked against the reference values (dashed lines) obtained by the HEOM method. For the case of $k_{\rm B}T=0.3\,\Gamma$, the PINN-DQME calculation is terminated at $t_{\rm end}=0.82\,\Gamma^{-1}$, as the minimization of loss function becomes increasingly demanding; see the main text for details. The MLP hyperparameters are chosen as $K=4$ and $N_{\rm h}=35$. For the case of $k_{{}_{\rm B}}T=3.0\,\Gamma$, the mapping functions $f_{1}(t)=t$ and $f_{2}(t)=f_{3}(t)=0$ are used in all subdomains. For the case of $k_{{}_{\rm B}}T=0.3\,\Gamma$, the mapping functions $f_{1}(t)=t$, $f_{2}(t)=t^{2}$, and $f_{3}(t)=t^{3}$ are used in the first two subdomains, and $f_{1}(t)=t$, $f_{2}(t)=t^{1.5}$, and $f_{3}(t)=t^{0.5}/(t+0.015\Gamma^{-1})$ are used in the later subdomains. System energetic parameters are chosen as (in units of $\Gamma$): $\epsilon_{0}=U_{0}/2=2$, $\Delta\epsilon=-7$, and $\Delta U=6$.
@@
Figure 3(a) and (b) exhibit the time-dependent current through the right reservoir $I_{\rm R}(t)$ and the occupation number of spin-up electrons on the impurity level $n_{\uparrow}(t)$, respectively, at high ($k_{\rm B}T=3.0\,\Gamma$) and low ($k_{\rm B}T=0.3\,\Gamma$) temperatures. The time-domain decomposition is illustrated through the dots on the lines. Within each subdomain, the time evolution of RDT is quantified by an MLP with three hidden layers ($K=4$), with each hidden layer comprised of 35 nodes ($N_{\rm h}=35$). The residual points (incremental time steps) are taken at intervals of $\Delta\tau\approx 0.015\,\Gamma^{-1}$. In the PINN-DQME framework, the expectation value of an observable $\hat{X}$, $X=\langle\hat{X}\rangle$, is evaluated from the parametrized RDT $\bm{\rho}_{\theta}$.
@@
For the case of $k_{\rm B}T=3.0\,\Gamma$, the PINN-DQME method yields highly accurate results that agree excellently with the reference data throughout the entire time domain, as evidenced by the rather small relative time-integrated errors, i.e., $\mathcal{E}_{X}<0.8\%$ with $X=I_{\rm R}$ and $n_{\uparrow}$, where the relative errors are defined by
@@
Here, $X^{{{\rm ref}}}$ denotes the reference values calculated by the HEOM method, and $t_{\rm end}$ is a chosen end point for time evolution. Particularly, the MLP trained within a subdomain $1.44\,\Gamma^{-1},1.67\,\Gamma^{-1}]$ can be straightforwardly extended to quantify the quantum dissipative dynamics up to the end time of $t_{\rm end}=2.3\,\Gamma^{-1}$ without changing any of its parameters. This highlights the remarkable practicality of PINN for capturing open quantum dynamics where non-Markovian effects are less prominent, because the dissipatons’ lifetimes are shorter at a higher temperature Li et al. (2023, 2024a).
@@
Figure 4: Comparison of electric current for two loss functions for the case of $k_{\rm B}T=0.3\,\Gamma$. The main panel shows the time evolution of the electric current obtained with the original loss function (blue line) and the modified loss function (orange line); see the main text for details. Purple and red dots mark the domain‑decomposition points $t_{p}$. The inset provides a zoomed‑in view of the third subdomain $0.3\,\Gamma^{-1},0.4\,\Gamma^{-1}]$, where the difference between the two loss functions becomes more apparent.
The considerable deviation observed in the low temperature case is a manifestation of error accumulation during the time evolution, which originates from the MLP’s inability to fully represent the quantitative non-Markovian memory effects within a given time subdomain. To elaborate, we modify the loss function in Eq. (18) that encodes the initial condition to $L_{{\rm I}_{p}}=\|\bm{\rho}_{\theta_{p}}(t_{p-1})-{\bm{\rho}}^{{\rm ref}}(t_{p-1})\|_{M}^{2}$ for the third subdomain $0.3\,\Gamma^{-1},0.4\,\Gamma^{-1}]$. Here, the reference values derived from the HEOM calculation ${\bm{\rho}}^{{\rm ref}}(t_{p-1})$ replaces the MLP-approximated counterpart $\bm{\rho}_{\theta_{p-1}}(t_{p-1})$. As demonstrated in Fig. 4, this adjustment yields substantially improved accuracy in the third subdomain. The result underscores the importance of maintaining high accuracy at subdomain boundaries. However, such a requirement can be highly demanding given the limited representational power of the MLP adopted in this study.
@@ III.2 Enhancing the performance of PINN-DQME
Next, we investigate how the choice of the feature mapping functions $\vec{f}(t)$ for the time nodes in the input layer affects the performance of PINN-DQME. Figure 6 compares results obtained using two different mappings within the first subdomain $0,0.228\,\Gamma^{-1}]$ for the case of $k_{\rm B}T=0.3\,\Gamma$: a simple linear mapping $\vec{f}_{a}(t):\{t\}$, and a richer polynomial mapping $\vec{f}_{b}(t):\{t,t^{2},t^{3}\}$. As depicted in Fig. 6(a), the richer feature set $\vec{f}_{b}(t)$ yields significantly more accurate population dynamics than the linear mapping $\vec{f}_{a}(t)$. This improvement is corroborated by the loss curves in Fig. 6(b), where $\vec{f}_{b}(t)$ achieves a lower final loss, confirming that enhanced input features facilitate both faster convergence and higher accuracy.
We have also explored several other techniques to enhance PINN‑DQME performance, such as replacing the simple MLP with a ResNet architecture He et al. (2016); Cheng and Zhang (2021) and modifying the initial‑condition loss to incorporate multiple residual points at subdomain boundaries. While these modifications yield modest accuracy improvements, they remain insufficient to reliably capture the full time evolution at $k_{\rm B}T=0.3\,\Gamma$. These results highlight the inherent limitations of the current PINN framework in representing complex non‑Markovian open quantum dynamics.
@@ IV Concluding remarks
The precise and efficient characterization of non-Markovian memory effects is often regarded as an “ultimate toughness test” Schuld et al. (2019) for numerical simulations of open quantum dynamics. Compared to recent advances in pure NQS approaches Cao et al. (2025), the current implementation of PINN-DQME adopts a more redical architecture: instead of representing the wave function at a single time point, a single neural network is optimized to represent the temporal evolution over a finite time interval. This approach offers potential gains in both speed and fidelity. However, this comes at the cost of a significantly more complex global optimization problem, necessitating the use of the accurate but computationally intensive BFGS algorithm and leading to reduced precision when simulating intricate non-Markovian dynamics. Nevertheless, its success in the high-temperature regime suggests that this direction is worthy of further exploration. The accuracy shortfall primarily stems from error accumulation in the time evolution, while the speed disadvantage is related to the use of a relatively slow BFGS optimizer. Future work may address these aspects by employing more powerful network architectures, such as transformers, or by incorporating more efficient training schemes.
@@ Acknowledgements.
Support from the National Natural Science Foundation of China (Grants No. 22393912, No. 22425301, No. 22373091 and No. 22573099), the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0303301 and No. 2021ZD0303306), the AI for Science Foundation of Fudan University (Grant No. FudanX24AI023), and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0450101) is gratefully acknowledged.

### 29 Apr 2026 v2

@@ III.1 Simulation of non-Markovian dissipative dynamics of Anderson impurity model
We consider an OQS comprising a localized impurity symmetrically coupled to two noninteracting electron reservoirs. This setup is important for understanding electron transport through molecular junctions. The OQS is described by the single-impurity Anderson model And6141 . The impurity Hamiltonian is $H_{{}_{\rm S}}(t)=\epsilon_{0}(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+U_{0}\hat{n}_{\uparrow}\hat{n}_{\downarrow}+{\Theta}(t-t_{0})[\Delta\epsilon(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+\Delta U\hat{n}_{\uparrow}\hat{n}_{\downarrow}$], where $\hat{n}_{s}$ is the occupation operator for spin-$s$ electrons, $\epsilon_{0}$ is the impurity energy, and $U_{0}$ is the Coulomb interaction energy. At time $t_{0}=0$, a sudden change shifts the impurity’s parameters by $\Delta\epsilon$ and $\Delta U$, and a bias voltage is simultaneously switched on, establishing a chemical potential difference between the reservoirs. The inset of Fig. 3(a) depicts the resulting open quantum dynamics. The impurity level shift triggers electron transfer from the reservoirs to the impurity on a relatively short timescale. Eventually a steady electric current flows through the impurity. At sufficiently low reservoir temperatures, environmental memory effects influence the electron transition process, resulting in non-Markovian oscillations. In the following, we also calculate the non-Markovian open quantum dynamics using the fermionic HEOM method implemented in the HEOM-QUICK program ye2016heom ; zhang2024heom , and the numerical outcomes are served as reference values for assessing the accuracy of the PINN-DQME method.
@@
Figure 3: (a) The evolution of the electric current flowing into the impurity from the right reservoir at different temperatures. Inset: schematic of quantum dissipative dynamics for an impurity coupled to left (L) and right (R) reservoirs with chemical potentials $\mu_{\rm L}$ and $\mu_{\rm R}$. (b) The evolution of the occupation number of the up spin electron in the impurity at two different temperatures. The purple and red dots represent the decomposition points $\{t_{p}\}$ for the time domain. The results of PINN-DQME (solid lines) are benchmarked against the reference values (dashed lines) obtained by the HEOM method. For the case of $k_{\rm B}T=0.3\,\Gamma$, the PINN-DQME calculation is terminated at $t_{\rm end}=0.82\,\Gamma^{-1}$, as the minimization of loss function becomes increasingly demanding; see the main text for details. The MLP hyperparameters are chosen as $K=4$ and $N_{\rm h}=35$. For the case of $k_{{}_{\rm B}}T=3.0\,\Gamma$, the mapping functions $f_{1}(t)=t$ and $f_{2}(t)=f_{3}(t)=0$ are used in all subdomains. For the case of $k_{{}_{\rm B}}T=2.0\,\Gamma$, the mapping functions $f_{1}(t)=t$, $f_{2}(t)=t^{2}$, and $f_{3}(t)=t^{3}$ are used in the first three subdomains, and $f_{1}(t)=t$, $f_{2}(t)=t^{1.5}$, and $f_{3}(t)=t^{0.5}/(t+0.015\Gamma^{-1})$ are used in the later subdomains. For the case of $k_{{}_{\rm B}}T=0.3\,\Gamma$, the mapping functions $f_{1}(t)=t$, $f_{2}(t)=t^{2}$, and $f_{3}(t)=t^{3}$ are used in the first two subdomains, and $f_{1}(t)=t$, $f_{2}(t)=t^{1.5}$, and $f_{3}(t)=t^{0.5}/(t+0.015\Gamma^{-1})$ are used in the later subdomains. System energetic parameters are chosen as (in units of $\Gamma$): $\epsilon_{0}=U_{0}/2=2$, $\Delta\epsilon=-7$, and $\Delta U=6$.
@@
Figure 3(a) and (b) exhibit the time-dependent current through the right reservoir $I_{\rm R}(t)$ and the occupation number of spin-up electrons on the impurity level $n_{\uparrow}(t)$, respectively, at high ($k_{\rm B}T=3.0\,\Gamma$), intermediate ($k_{\rm B}T=2.0\,\Gamma$) and low ($k_{\rm B}T=0.3\,\Gamma$) temperatures. The time-domain decomposition is illustrated through the dots on the lines. Within each subdomain, the time evolution of RDT is quantified by an MLP with three hidden layers ($K=4$), with each hidden layer comprised of 35 nodes ($N_{\rm h}=35$). The residual points (incremental time steps) are taken at intervals of $\Delta\tau\approx 0.015\,\Gamma^{-1}$. In the PINN-DQME framework, the expectation value of an observable $\hat{X}$, $X=\langle\hat{X}\rangle$, is evaluated from the parametrized RDT $\bm{\rho}_{\theta}$.
@@
For the cases of $k_{\rm B}T=3.0\,\Gamma$ and $k_{\rm B}T=2.0\,\Gamma$, the PINN-DQME method yields highly accurate results that agree excellently with the reference data throughout the entire time domain, as evidenced by the rather small relative time-integrated errors, i.e., $\mathcal{E}_{X}<1.7\%$ with $X=I_{\rm R}$ and $n_{\uparrow}$, where the relative errors are defined by
@@
Here, $X^{{{\rm ref}}}$ denotes the reference values calculated by the HEOM method, and $t_{\rm end}$ is a chosen end point for time evolution. Particularly, the MLP trained within a subdomain $1.44\,\Gamma^{-1},1.67\,\Gamma^{-1}]$ at $k_{\rm B}T=3.0\,\Gamma$ can be straightforwardly extended to quantify the quantum dissipative dynamics up to the end time of $t_{\rm end}=2.3\,\Gamma^{-1}$ without changing any of its parameters. This highlights the remarkable practicality of PINN for capturing open quantum dynamics where non-Markovian effects are less prominent, because the dissipatons’ lifetimes are shorter at higher temperatures li2023dissipatons ; li2024quantum .
@@
Figure 4: (a) Comparison of electric current for two loss functions for the case of $k_{\rm B}T=0.3\,\Gamma$. The main panel shows the time evolution of the electric current obtained with the original loss function (blue line) and the modified loss function (orange line); see the main text for details. Purple and red dots mark the domain‑decomposition points $t_{p}$. The inset provides a zoomed‑in view of the third subdomain $0.3\,\Gamma^{-1},0.4\,\Gamma^{-1}]$, where the difference between the two loss functions becomes more apparent. (b) Imaginary part of the hybrid correlation function $C^{+}(t)$ at different temperatures. The dashed horizontal line is at $y=10^{-1}\,\Gamma^{2}$.
The considerable deviation observed in the low temperature case is a manifestation of error accumulation during the time evolution, which originates from the MLP’s inability to fully represent the quantitative non-Markovian memory effects within a given time subdomain. To elaborate, we modify the loss function in Eq. (18) that encodes the initial condition to $L_{{\rm I}_{p}}=\|\bm{\rho}_{\theta_{p}}(t_{p-1})-{\bm{\rho}}^{{\rm ref}}(t_{p-1})\|_{M}^{2}$ for the third subdomain $0.3\,\Gamma^{-1},0.4\,\Gamma^{-1}]$. Here, the reference values derived from the HEOM calculation ${\bm{\rho}}^{{\rm ref}}(t_{p-1})$ replaces the MLP-approximated counterpart $\bm{\rho}_{\theta_{p-1}}(t_{p-1})$. As demonstrated in Fig. 4(a), this adjustment yields substantially improved accuracy in the third subdomain. The result underscores the importance of maintaining high accuracy at subdomain boundaries. However, such a requirement can be highly demanding given the limited representational power of the MLP adopted in this study.
This error accumulation primarily stems from the strongly non‑Markovian memory at low temperatures. Non‑Markovianity describes the backflow of information from the environment to the system. Within the PINN framework, this backflow leads to complex time-dependent evolution that is more prone to error accumulation. Consequently, when the environment exhibits stronger non‑Markovian memory, initial errors tend to persist and propagate over longer time intervals, exerting a prolonged influence on the subsequent temporal evolution. Such error accumulation is particularly detrimental to PINN‑based simulations. Figure 4(b) illustrates the imaginary part of the hybrid correlation function $C^{+}(t)$ at different temperatures. The dashed horizontal line marks $y=10^{-1}\,\Gamma^{2}$. The correlation function at $k_{\rm B}T=0.3\,\Gamma$ decays significantly more slowly than that at $k_{\rm B}T=3.0\,\Gamma$, indicating stronger non‑Markovian effects at lower temperatures.
Table 1: Comparison of parameter counts for the PINN-DQME and NQS-DQME methods.
$k_{\rm B}T/\Gamma$ $3.0$ $2.0$ $0.3$
$N_{\rm para\,(PINN)}$ per subdomain 6182 8472 10152
$N_{\rm para\,(NQS)}$ per time point 3712 4584 10510
Then, we compare the computational cost of the PINN-DQME method and the NQS-DQME method cao2026simulating . We first compare parameter counts in Table 1. PINN‑DQME uses similar parameters per subdomain as NQS‑DQME per time point, but requires only $\sim$10 subdomains for the full evolution versus $\sim$1000 time points for NQS‑DQME, resulting in a much smaller total. A direct fair runtime comparison is challenging because the two approaches are implemented in different frameworks, and their convergence criteria, architectures, and hyperparameters differ significantly, which could make a simple runtime comparison misleading. Nevertheless, a partial quantitative comparison is still possible: the time for PINN to reduce the loss from a suitable initialization to $10^{-4}$ is of the same order as the time for NQS to evolve through the same subdomain. This suggests that PINN‑DQME has the potential to outperform NQS‑DQME in weakly non‑Markovian regimes where only moderate precision is required.
@@ III.2 Enhancing the performance of PINN-DQME
Next, we investigate how the choice of the feature mapping functions $\vec{f}(t)$ for the time nodes in the input layer affects the performance of PINN-DQME. Figure 6 compares results obtained using two different mappings within the first subdomain $0,0.228\,\Gamma^{-1}]$ for the case of $k_{\rm B}T=0.3\,\Gamma$: a simple linear mapping $\vec{f}_{a}(t):\{t\}$, and a richer polynomial mapping $\vec{f}_{b}(t):\{t,t^{2},t^{3}\}$. As depicted in Fig. 6(a), the richer feature set $\vec{f}_{b}(t)$ yields significantly more accurate population dynamics than the linear mapping $\vec{f}_{a}(t)$. This improvement is corroborated by the loss curves in Fig. 6(b), where $\vec{f}_{b}(t)$ achieves a lower final loss, confirming that enhanced input features facilitate both faster convergence and higher accuracy. Here, the time-feature mapping functions were chosen based on simplicity and the physical behavior of the system. Specifically, $\vec{f}_{a}(t)=\{t\}$ and the richer polynomial mapping $\vec{f}_{b}(t)=\{t,t^{2},t^{3}\}$ were selected for their simplicity, although other mappings are also possible. However, in our tests, exponential and logarithmic mappings (e.g. $\{e^{t},\ \log t\}$) did not perform as well as the polynomial ones. For later subdomains, the term $t^{0.5}/(t+0.015\Gamma^{-1})$ was designed to capture the decreasing trend of the current. Its numerator $t^{0.5}$ ensures that the term vanishes at $t=0$, similar to the polynomial mappings, while the constant $0.015\,\Gamma^{-1}$ in the denominator regularizes the expression, preventing divergence and ensuring a finite value at $t=0$.
We have also explored several other techniques to enhance PINN‑DQME performance, such as replacing the simple MLP with a ResNet architecture he2016deep ; cheng2021deep and modifying the initial‑condition loss to incorporate multiple residual points at subdomain boundaries. While these modifications yield modest accuracy improvements, they remain insufficient to reliably capture the full time evolution at $k_{\rm B}T=0.3\,\Gamma$. These results highlight the inherent limitations of the current PINN framework in representing complex non‑Markovian open quantum dynamics.
@@ IV Concluding remarks
The precise and efficient characterization of non-Markovian memory effects is often regarded as an “ultimate toughness test” Schuld19 for numerical simulations of open quantum dynamics. Compared to recent advances in pure NQS approaches cao2026simulating , the current implementation of PINN-DQME adopts a more radical architecture: instead of representing the wave function at a single time point, a single neural network is optimized to represent the temporal evolution over a finite time interval. This approach offers potential gains in both speed and fidelity. However, this comes at the cost of a significantly more complex global optimization problem, necessitating the use of the accurate but computationally intensive BFGS algorithm and leading to reduced precision when simulating intricate non-Markovian dynamics. Nevertheless, its success in the high-temperature regime suggests that this direction is worthy of further exploration. The accuracy shortfall primarily stems from error accumulation in the time evolution, while the speed disadvantage is related to the use of a relatively slow BFGS optimizer. Future work may address these aspects by employing more powerful network architectures, such as transformers, or by incorporating more efficient training schemes.
@@ Acknowledgements.
Support from the National Natural Science Foundation of China (Grants No. 22393912, No. 22425301, No. 22373091 and No. 22573099), the Quantum Science and Technology–National Science and Technology Major Project (Grants No. 2021ZD0303301 and No. 2021ZD0303306), the AI for Science Foundation of Fudan University (Grant No. FudanX24AI023), and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0450101) is gratefully acknowledged.
@@
Figure 7: Comparison of the electric current at different temperatures in the first subdomain $0,0.228\,\Gamma^{-1}]$ with a similar loss function value ($\sim 3\times 10^{-4}$). The mapping functions are $f_{1}(t)=t$, $f_{2}(t)=t^{2}$, and $f_{3}(t)=t^{3}$ for all cases. Dots on each curve mark the residual points used during optimization of PINN.
Appendix A Exploring the reliable temperature regime of PINN-DQME method
Figure 7 compares the electric current at different temperatures in the first subdomain $0,0.228\,\Gamma^{-1}]$ while maintaining a similar loss function value ($\sim 3\times 10^{-4}$). It is observed that, as the temperature decreases, the deviation from reference values increases significantly, despite the comparable loss. This growing deviation reflects the heightened precision requirement at lower temperatures due to the enhanced non-Markovianity.
In our numerical tests, we find that reducing the loss function below $10^{-4}$ is difficult to achieve. As demonstrated in Fig. 7, the precision required for temperatures below $\Gamma$ is less than $10^{-4}$. Therefore, the reliable temperature regime of the present PINN-DQME method is $k_{\rm B}T>\Gamma$.
Appendix B Comparison of different optimizers
Figure 8: Comparison of different optimizers at the temperature $k_{\rm B}T=2.0\,\Gamma$ in the first subdomain $0,0.228\,\Gamma^{-1}]$. (a) Training loss (log scale) as a function of optimization steps. (b) Electric current obtained by different optimizers at the final points of (a). The mapping functions are $f_{1}(t)=t$, $f_{2}(t)=t^{2}$, and $f_{3}(t)=t^{3}$ for all cases. Dots on each curve mark the residual points used during optimization of PINN.
Figure 8 compares the results of two common optimizers in PINN (Adam kingma2014adam and L-BFGS-B zhu1997algorithm ) with that of the BFGS method. Figure 8(a) shows the decreasing process of the training loss of different optimizers, illustrating that while BFGS reduces the loss to $10^{-4}$, the other two methods only achieve a loss of $10^{-3}$. The L‑BFGS‑B method terminates early because the relative loss reduction fell below the convergence threshold. In contrast, the Adam method does not employ such a convergence criterion; instead, it continues training with a fixed learning rate of $4\times 10^{-7}$. Figure 8(b) compares the final results obtained by three optimizers. The results from the L-BFGS-B and Adam methods deviate significantly from the reference values, which will lead to large errors in subsequent evolution. Only the BFGS method yeiled the results that agree with the reference values, further confirming that the BFGS method is required for achieving sufficient accuracy.